This section contains 5 questions. The answer to each of the questions is a single digit integer ranging from 0 to 9. The error in the measurement of volume of sphere is 1.5%. The percentage error in the measurement of its diameter is A × 10–1. Find the value of A.

Grade:upto college level

2 Answers

Radhika Batra
10 years ago

5
V = 4/3 ? r3 = ?/6 d2
% error in volume = 3 × % error in diameter.
1.5 = 3 × % error in diameter = 0.5% = (5 × 10–1)%
A = 5

Chainsingh
6 years ago
4/3πr³ is volume in radius If we take it in diameter than it becomesπ\6d³π/6 is constant Therefore error is 3*d=1.5So error come is 0.5